OpenStudy (anonymous):

1/ sqr root 3t^4 answer should be 3^-1/2 t^-2 but I got 1/ 3^1/2 t^2 so would I just flip it ?

OpenStudy (campbell_st):

so are you trying to simplify \[\frac{1}{\sqrt{3t^4}}\]

OpenStudy (campbell_st):

ok... you're writing it in index form so the index for roots or radicals is \[x^{\frac{1}{2}} = \sqrt {x}\] so you have \[\frac{1}{(3t^4)^{\frac{1}{2}}}\] fractions have negative indices \[x^{-a} = \frac{1}{x^a}\] so you have \[(3t^4)^{-\frac{1}{2}}\] lastly use the power of a power rule \[(x^ay^b)^c = x^{a \times c} y^{b \times c} \] that will give the answer you need

OpenStudy (anonymous):

waiiittttt, where did the negative come from? :S

OpenStudy (campbell_st):

terms in denominators can be written with negative indices